Please explain how to do with a Business Calculator

A 8-year annuity of 16 $9,300 semiannual payments will begin 10.5 years from now, with the first payment coming 11 years from now.

Requirement 1:

If the discount rate is 12 percent compounded semiannually, what is the value of this annuity eight years and six years from now? (Enter rounded answers as directed, but do not use rounded numbers in intermediate calculations. Round your answers to 2 decimal places (e.g., 32.16).)

Value of the annuity
  Eight years $          
  Six years $          
Requirement 2:

What is the value of the annuity today?
Answer #1

Value of Annuity 10.5 years from now,

PV of Annuity = P[(1 - (1+r)-n)/r]

Value of Annuity at 10.5 years = $93,984.83

a.

Value of Annuity at the end of Year 8 = 93,984.83/(1.06)5

Value of Annuity at the end of Year 8 = $70,230.93

b.

Value of Annuity at the end of Year 6 = 93,984.83/(1.06)9

Value of Annuity at the end of Year 6 = $55,629.48

c.

Value of Annuity Today = 93,984.83/(1.06)21

Value of Annuity Today = $27,646.15